OpenStudy (anonymous):

f(1) = 35, f(n) = f(1) + f(n-1) for n > 1 Use the recursive formula to make a table of values for 1 ≤ n ≤ 5. Show your calculations. Explain what your table means.

OpenStudy (michele_laino):

please note that: \[f(2)=f(1)+f(1)=2f(1)\] \[f(3)=f(1)+f(2)=f(1)+2f(1)=3f(1)\] so I congecture this: \[f(n)=nf(1)(*)\] and I suppose it is valid, now for n+1, we have: \[f(n+1)=f(1)+f(n)=f(1)+nf(1)=(n+1)f(n)\] so (*) it is valid for all n such that: \[\forall n \in \mathbb{N},n \ge1\]

OpenStudy (anonymous):

Okay thanks.

OpenStudy (michele_laino):

thanks!

OpenStudy (michele_laino):

f(1)=35; f(2)=2*f(1)=70 please continue

OpenStudy (michele_laino):

f(3)=3*f(1)=....